# The numeraire e-variable and reverse information projection

### v3

@@ 4 Finding the numeraire
If ${\mathcal{P}}$ is convex, $H({\mathsf{Q}}\mid p)<\infty$ for some $p\in{\mathcal{P}}$, and $(p_{n})_{n\in{\mathbb{N}}}$ is an entropy minimizing sequence in ${\mathcal{P}}$, meaning that $H({\mathsf{Q}}\mid p_{n})\to\inf_{p\in{\mathcal{P}}}H({\mathsf{Q}}\mid p)$, then $\log p_{n}$ converges to $\log p^{*}$ in $L^{1}({\mathsf{Q}})$, where $p^{*}$ is the RIPr.
@@
Since the quantities involved are finite for $n$ large enough, $\lim_{n\to\infty}H({\mathsf{Q}}\mid p_{n}\leadsto{\mathcal{P}})=\lim_{n\to\infty}H({\mathsf{Q}}\mid p_{n})-\inf_{p\in{\mathcal{P}}}H({\mathsf{Q}}\mid p)=0$, and the result follows from Corollary 4.9. ∎
@@ 5.1 A parametric null
We start with a simple parametric example from Lardy et al. (2023) involving a real-valued observation $Z$ in ${\mathcal{Z}}={\mathbb{R}}$, and then generalize our observations at the end of this subsection. The null hypothesis ${\mathcal{P}}$ consists of two unit-variance Gaussians ${\mathsf{P}}_{1}$ and ${\mathsf{P}}_{2}$ with mean $+1$ and $-1$, and the alternative hypothesis ${\mathsf{Q}}$ is the standard Cauchy distribution. We claim that the RIPr is ${\mathsf{P}}^{*}=({\mathsf{P}}_{1}+{\mathsf{P}}_{2})/2$. This probability measure is equivalent to ${\mathsf{Q}}$ and belongs to ${\mathcal{P}}^{\circ\circ}$, which always contains the convex hull of ${\mathcal{P}}$. Furthermore, writing $p_{1},p_{2},p^{*},q$ for the Lebesgue densities, we have $p_{2}(z)=p_{1}(-z)$ and $q(z)=q(-z)$ and hence
@@ 5.2 Bounded mean
To find a candidate numeraire, we observe that two natural e-variables are $Z/\mu$ and the constant one. All convex combinations of these are also e-variables; equivalently, $1+\lambda(Z-\mu)$ is an e-variable for each $\lambda\in[0,\mu^{-1}]$. We now look for a log-optimal e-variable in this class by directly maximizing $f(\lambda)={\mathbb{E}}_{\mathsf{Q}}[\log(1+\lambda(Z-\mu))]$ over $\lambda\in[0,\mu^{-1}]$. This is a strictly concave function whose derivative $f^{\prime}(\lambda)={\mathbb{E}}_{\mathsf{Q}}[(Z-\mu)/(1+\lambda(Z-\mu))]$ satisfies $f^{\prime}(0)=\frac{1}{2}-\mu>0$ and $f^{\prime}(\mu^{-1})={\mathbb{E}}_{\mathsf{Q}}[\mu-\mu^{2}/Z]=-\infty$. Thus there is a unique interior maximizer $\lambda^{*}\in(0,\mu^{-1})$ which is characterized by the first-order condition
@@
Taking $\lambda=0$ and $\lambda=1/\mu$ we see that (4.2) of Corollary 4.2 is satisfied and, hence, that $X^{*}$ is the numeraire.
@@ 5.4 Symmetric distributions
We claim that the set of all e-variables is
@@
where we recall that a function $\phi$ is odd if $\phi(z)+\phi(-z)=0$ for all $z\in{\mathbb{R}}$. Indeed, any $X$ of this form satisfies ${\mathbb{E}}_{\mathsf{P}}[X]\leq 1+{\mathbb{E}}_{\mathsf{P}}[\phi(Z)]=1$, where the symmetry of $Z$ and the oddness of $\phi$ were used to get ${\mathbb{E}}_{\mathsf{P}}[\phi(Z)]={\mathbb{E}}_{\mathsf{P}}[\phi(-Z)]=-{\mathbb{E}}_{\mathsf{P}}[\phi(Z)]$ and hence ${\mathbb{E}}_{\mathsf{P}}[\phi(Z)]=0$. Conversely, for any e-variable $X=f(Z)$ we may write $X=\frac{1}{2}(f(Z)+f(-Z))+\phi(Z)$ where $\phi(z)=\frac{1}{2}(f(z)-f(-z))$ is the odd part of $f(z)$. For any $z\in{\mathbb{R}}$ we use the symmetric distribution ${\mathsf{P}}=\frac{1}{2}(\delta_{z}+\delta_{-z})$ and the fact that $X$ is an e-variable to get $\frac{1}{2}(f(z)+f(-z))={\mathbb{E}}_{\mathsf{P}}[f(Z)]\leq 1$. Hence $X\leq 1+\phi(Z)$ as required. We could also have reached this conclusion by using that every admissible e-variable is of the form $1+\phi(Z)$ for some odd function $\phi$ as shown by Ramdas et al. (2020).
@@ 5.5 Exponential family with one-dimensional sufficient statistic
Here ${\mathcal{Z}}$ can be general and is equipped with a reference measure $\mu$. Consider an exponential family of densities
@@ 6 Beyond the logarithm: reverse Rényi projection
Thanks to (6.1), ${\mathsf{P}}^{*}_{\gamma}$ belongs to ${\mathcal{P}}^{\circ\circ}$. In analogy with (3.3) we expect ${\mathsf{P}}^{*}_{\gamma}$ to minimize a suitable dual objective function. To discover its form, we look for a weak duality inequality and introduce the Legendre transform $V(y)=\sup_{x>0}\{U(x)-xy\}$. Evaluating the supremum shows that
@@
Now, for any e-variable $X$, any ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$, and any $y\geq 0$, we have
@@
To obtain a tight bound we minimize the right-hand side over $y$. A computation shows that the minimizer is $y={\mathbb{E}}_{\mathsf{Q}}[(d{\mathsf{P}}^{a}/d{\mathsf{Q}})^{1-1/\gamma}]^{\gamma}$, which is finite, possibly zero, thanks to Jensen’s inequality. With this choice we get, for any e-variable $X$ and any ${\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}$, the weak duality inequality
@@
$\inf_{{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}}\frac{1}{1-\gamma}{\mathbb{E}}_{\mathsf{Q}}\left[\left(\frac{d{\mathsf{P}}^{a}}{d{\mathsf{Q}}}\right)^{1-1/\gamma}\right]^{\gamma}$
@@
Thus ${\mathsf{P}}^{*}_{\gamma}$ also minimizes the Rényi divergence, and as such it is known as the reverse Rényi projection of order $1/\gamma$ of ${\mathsf{Q}}$ on ${\mathcal{P}}^{\circ\circ}$. In particular, we see that the reverse Rényi projection always exists if the order $1/\gamma$ is in $(0,1)$. We summarize the above observations, along with a few others, in the following theorem.
@@
$\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{\mathsf{Q}}[U(X)]=\frac{1}{1-\gamma}\exp\left((1-\gamma)\inf_{{\mathsf{P}}\in{\mathcal{P}}^{\circ\circ}}R_{1/\gamma}({\mathsf{Q}}\mid{\mathsf{P}})\right),$ (6.3)
and the supremum and infimum are attained by $X^{*}_{\gamma}\in{\mathcal{E}}$ and ${\mathsf{P}}^{*}_{\gamma}\in{\mathcal{P}}^{\circ\circ}$ related by (6.2). Furthermore, $X^{*}_{\gamma}$ is the ${\mathsf{Q}}$-almost surely unique maximizer, and ${\mathsf{P}}^{*}_{\gamma}$ is the unique minimizer among the elements of ${\mathcal{P}}^{\circ\circ}$ that are absolutely continuous with respect to ${\mathsf{Q}}$. Finally, $X^{*}_{\gamma}$ can be expressed in terms of ${\mathsf{P}}^{*}_{\gamma}$ by
@@
The equality in (6.3) and the form of the solutions were shown above. The ${\mathsf{Q}}$-almost sure uniqueness of $X^{*}_{\gamma}$ follows from strict concavity of the function $f(X)={\mathbb{E}}_{\mathsf{Q}}[U(X)]$ on ${\mathcal{E}}$. Since $\gamma>1$, the Rényi divergence is a decreasing transformation of the function $g({\mathsf{P}})={\mathbb{E}}_{\mathsf{Q}}[(d{\mathsf{P}}^{a}/d{\mathsf{Q}})^{1-1/\gamma}]$, which is thus maximized by ${\mathsf{P}}^{*}_{\gamma}$. The function $g({\mathsf{P}})$ is strictly concave on the set of elements of ${\mathcal{P}}^{\circ\circ}$ that are absolutely continuous with respect to ${\mathsf{Q}}$. Thus ${\mathsf{P}}^{*}_{\gamma}$ is the unique maximizer of $g({\mathsf{P}})$, and the unique minimizer of the Rényi divergence. Finally, (6.4) is obtained by noting that
@@
Lardy et al. (2023, Section V) mention the e-variable $(d{\mathsf{P}}^{*}_{\gamma}/d{\mathsf{Q}})^{-1/\gamma}$ constructed using the reverse Rényi projection. This e-variable is strictly smaller than $X^{*}_{\gamma}$, because the denominator in (6.4) is strictly smaller than one by Jensen’s inequality.
@@
7 Composite alternatives and universal inference
We now briefly discuss implications of our work for composite alternatives, as well as its relation to universal inference (Wasserman et al., 2020). We purposefully keep the discussion informal.
Method of mixtures and plug-in method.
Despite our focus on point alternatives, there are two natural ways in which this theory has immediate implications for composite alternatives ${\mathcal{Q}}$. When testing ${\mathcal{P}}$ against ${\mathcal{Q}}$ for i.i.d. data $(Z_{t})_{t\in{\mathbb{N}}}$, one sensible goal is to derive e-variables $X_{t}$ depending on the sample $Z_{1},\ldots,Z_{t}$ that are ‘asymptotically optimal’ in the sense that for all ${\mathsf{Q}}\in{\mathcal{Q}}$, we would have ${\mathbb{E}}_{{\mathsf{Q}}^{\otimes t}}[t^{-1}\log X_{t}]\to H({\mathsf{Q}}\mid{\mathsf{P}}^{*}({\mathsf{Q}}))$ as the sample size $t$ increases to infinity, where we use ${\mathsf{P}}^{*}({\mathsf{Q}})$ to denote the RIPr of ${\mathsf{Q}}$ onto ${\mathcal{P}}^{\circ\circ}$. Because $H({\mathsf{Q}}\mid{\mathsf{P}}^{*}({\mathsf{Q}}))$ is the highest achievable growth rate if the data is generated according to ${\mathsf{Q}}$, the above says that this optimal growth rate is achieved asymptotically no matter which ${\mathsf{Q}}\in{\mathcal{Q}}$ is the true data generating distribution. This is clearly a desirable feature whenever the data comes from an unknown element of ${\mathcal{Q}}$.
The literature deals with composite alternatives in two ways (Ramdas et al., 2023). The first is called the method of mixtures: using a distribution $\pi$ over ${\mathcal{Q}}$ and assuming for simplicity that the elements of ${\mathcal{Q}}$ are mutually equivalent, one takes the mixture likelihood ratio $X_{t}=\int_{\mathcal{Q}}\prod_{i\leq t}\frac{d{\mathsf{Q}}}{d{\mathsf{P}}^{*}({\mathsf{Q}})}(Z_{i})\pi(d{\mathsf{Q}})$. The second method is called the plug-in method: using part of the sample $Z_{1},\ldots,Z_{t}$ one chooses a particular representative element $\hat{\mathsf{Q}}_{t}$ of ${\mathcal{Q}}$ and uses this to obtain the corresponding numeraire as a function of the remaining part of the sample. This method is an instance of sample-splitting, used for instance in universal inference (Wasserman et al., 2020). In either case we get valid e-variables, and under relatively weak conditions on $\pi$ or $\hat{\mathsf{Q}}_{t}$ one can prove that the resulting e-variable is indeed asymptotically optimal. Establishing tight necessary and sufficient conditions for such asymptotic optimality is left to future work.
Comparison to universal inference.

### v4

@@ 4.2 When a reference measure exists
Assume that ${\mathsf{Q}}\ll{\mathcal{P}}$ and that all elements of ${\mathcal{P}}$ are absolutely continuous with respect to a probability measure $\mu$. If ${\mathcal{P}}$ is convex, $H({\mathsf{Q}}\mid p)<\infty$ for some $p\in{\mathcal{P}}$, and $(p_{n})_{n\in{\mathbb{N}}}$ is an entropy minimizing sequence in ${\mathcal{P}}$, meaning that $H({\mathsf{Q}}\mid p_{n})\to\inf_{p\in{\mathcal{P}}}H({\mathsf{Q}}\mid p)$, then $\log p_{n}$ converges to $\log p^{*}$ in $L^{1}({\mathsf{Q}})$, where $p^{*}$ is the RIPr. In particular, if some $p^{\prime}\in{\mathcal{P}}$ achieves the infimum, then $p^{\prime}$ is the RIPr. Finally, $H(Q\mid p^{*})=\inf_{p\in{\mathcal{P}}}H({\mathsf{Q}}\mid p)$.
@@
Since the quantities involved are finite for $n$ large enough, $\lim_{n\to\infty}H({\mathsf{Q}}\mid p_{n}\leadsto{\mathcal{P}})=\lim_{n\to\infty}H({\mathsf{Q}}\mid p_{n})-\inf_{p\in{\mathcal{P}}}H({\mathsf{Q}}\mid p)=0$, and the first claim follows from Corollary 4.10. The second claim follows by simply taking $p_{n}=p^{\prime}$ for all $n$. The final claim follows because $H({\mathsf{Q}}\mid p_{n})\to H({\mathsf{Q}}\mid p^{*})$ thanks to the $L^{1}(Q)$ convergence of the log-densities. ∎
Corollaries 4.10 and 4.11 confirm that the definition of the RIPr used in this paper, Definition 3.4, is consistent with existing definitions in the literature. Specifically, Li (1999) defines the RIPr as the density $p^{*}$ with the property that whenever $(p_{n})_{n\in{\mathbb{N}}}$ is a sequence in ${\mathcal{P}}$ such that $H({\mathsf{Q}}\mid p_{n})\to\inf_{p\in{\mathcal{P}}}H({\mathsf{Q}}\mid p)$, we have $\log p_{n}\to\log p^{*}$ in $L^{1}({\mathsf{Q}})$. Corollary 4.11 states that this indeed yields the RIPr of Definition 3.4. Similarly, Lardy et al. (2024) define the RIPr as the density $p^{*}$ with the property that whenever $(p_{n})_{n\in{\mathbb{N}}}$ is a sequence in ${\mathcal{P}}$ such that $H({\mathsf{Q}}\mid p_{n}\leadsto{\mathcal{P}})\to\inf_{p\in{\mathcal{P}}}H({\mathsf{Q}}\mid p\leadsto{\mathcal{P}})$ (which must be zero whenever it is finite), $p_{n}$ converges to ${\mathsf{P}}^{*}$ in a certain metric which is equivalent to $L^{1}({\mathsf{Q}})$-convergence of the log-densities. Thanks to Corollary 4.10, this again yields the RIPr of Definition 3.4.
@@ 5.1 A parametric null
We start with a simple parametric example from Lardy et al. (2024) involving a real-valued observation $Z$ in ${\mathcal{Z}}={\mathbb{R}}$, and then generalize our observations at the end of this subsection. The null hypothesis ${\mathcal{P}}$ consists of two unit-variance Gaussians ${\mathsf{P}}_{1}$ and ${\mathsf{P}}_{2}$ with mean $+1$ and $-1$, and the alternative hypothesis ${\mathsf{Q}}$ is the standard Cauchy distribution. This example was considered by the aforementioned authors because $H(Q\mid P_{j})=\infty$ for $j\in\{1,2\}$.
We claim that the RIPr is ${\mathsf{P}}^{*}=({\mathsf{P}}_{1}+{\mathsf{P}}_{2})/2$. This probability measure is equivalent to ${\mathsf{Q}}$ and belongs to ${\mathcal{P}_{\textnormal{eff}}}$, which always contains the convex hull of ${\mathcal{P}}$. Furthermore, writing $p_{1},p_{2},p^{*},q$ for the Lebesgue densities, we have $p_{2}(z)=p_{1}(-z)$ and $q(z)=q(-z)$ and hence
@@ 5.2 Bounded mean
To find a candidate numeraire, we observe that two natural e-variables are $Z/\mu$ and the constant one. All convex combinations of these are also e-variables; equivalently, $1+\lambda(Z-\mu)$ is an e-variable for each $\lambda\in[0,\mu^{-1}]$. We now look for a log-optimal e-variable in this class by directly maximizing $f(\lambda)={\mathbb{E}}_{\mathsf{Q}}[\log(1+\lambda(Z-\mu))]$ over $\lambda\in[0,\mu^{-1}]$. This is a strictly concave function whose derivative $f^{\prime}(\lambda)={\mathbb{E}}_{\mathsf{Q}}[(Z-\mu)/(1+\lambda(Z-\mu))]$ satisfies $f^{\prime}(0)={\mathbb{E}}_{\mathsf{Q}}[Z]-\mu>0$ and $f^{\prime}(\mu^{-1})={\mathbb{E}}_{\mathsf{Q}}[\mu-\mu^{2}/Z]=-\infty$. Thus there is a unique interior maximizer $\lambda^{*}\in(0,\mu^{-1})$, which is characterized by the first-order condition
@@
Taking $\lambda=0$ and $\lambda=1/\mu$ we see that (4.2) of Corollary 4.2 is satisfied and, hence, that $X^{*}$ is the numeraire and the RIPr ${\mathsf{P}}^{*}$ belongs to ${\mathcal{P}}$.
@@ 5.4 Symmetric distributions
This expression also appears in Koning (2023, Section 5–6). We claim that the set of all e-variables is
@@
where we recall that a function $\phi$ is odd if $\phi(z)+\phi(-z)=0$ for all $z\in{\mathbb{R}}$. Indeed, any $X$ of this form satisfies ${\mathbb{E}}_{\mathsf{P}}[X]\leq 1+{\mathbb{E}}_{\mathsf{P}}[\phi(Z)]=1$, where the symmetry of $Z$ and the oddness of $\phi$ were used to get ${\mathbb{E}}_{\mathsf{P}}[\phi(Z)]={\mathbb{E}}_{\mathsf{P}}[\phi(-Z)]=-{\mathbb{E}}_{\mathsf{P}}[\phi(Z)]$ and hence ${\mathbb{E}}_{\mathsf{P}}[\phi(Z)]=0$. Conversely, for any e-variable $X=f(Z)$ we may write $X=\frac{1}{2}(f(Z)+f(-Z))+\phi(Z)$ where $\phi(z)=\frac{1}{2}(f(z)-f(-z))$ is the odd part of $f(z)$. For any $z\in{\mathbb{R}}$ we use the symmetric distribution ${\mathsf{P}}=\frac{1}{2}(\delta_{z}+\delta_{-z})$ and the fact that $X$ is an e-variable to get $\frac{1}{2}(f(z)+f(-z))={\mathbb{E}}_{\mathsf{P}}[f(Z)]\leq 1$. Hence $X\leq 1+\phi(Z)$ as required. This is a different conclusion from the fact that every admissible and exact e-variable is of the form $1+\phi(Z)$ for some odd function $\phi$, as shown by Ramdas et al. (2020).
@@ 5.5 Exponential family with one-dimensional sufficient statistic
Here ${\mathcal{Z}}$ can be general and is equipped with a reference measure $\mu$. The following example shows how to use the results in Section 4.2 to obtain the result in (Grünwald et al., 2024, Example 4). Consider an exponential family of densities
@@ 6 Beyond the logarithm: reverse Rényi projection
Thanks to (6.1), ${\mathsf{P}}^{*}_{\gamma}$ belongs to ${\mathcal{P}_{\textnormal{eff}}}$. In analogy with (3.3) we expect ${\mathsf{P}}^{*}_{\gamma}$ to minimize a suitable dual objective function. To discover its form, we look for a weak duality inequality and introduce the Legendre transform $V(y)=\sup_{x>0}\{U(x)-xy\}$. Evaluating the supremum shows that
@@
Now, for any e-variable $X$, any ${\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}$, and any $y\geq 0$, we have
@@
To obtain a tight bound we minimize the right-hand side over $y$. A computation shows that the minimizer is $y={\mathbb{E}}_{\mathsf{Q}}[(d{\mathsf{P}}^{a}/d{\mathsf{Q}})^{1-1/\gamma}]^{\gamma}$, which is finite, possibly zero, thanks to Jensen’s inequality. With this choice we get, for any e-variable $X$ and any ${\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}$, the weak duality inequality
@@
$\inf_{{\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}}\frac{1}{1-\gamma}{\mathbb{E}}_{\mathsf{Q}}\left[\left(\frac{d{\mathsf{P}}^{a}}{d{\mathsf{Q}}}\right)^{1-1/\gamma}\right]^{\gamma}$
@@
Thus ${\mathsf{P}}^{*}_{\gamma}$ also minimizes the Rényi divergence, and as such it is known as the reverse Rényi projection of order $1/\gamma$ of ${\mathsf{Q}}$ on ${\mathcal{P}_{\textnormal{eff}}}$. In particular, we see that the reverse Rényi projection always exists if the order $1/\gamma$ is in $(0,1)$. We summarize the above observations, along with a few others, in the following theorem.
@@
$\sup_{X\in{\mathcal{E}}}{\mathbb{E}}_{\mathsf{Q}}[U(X)]=\frac{1}{1-\gamma}\exp\left((1-\gamma)\inf_{{\mathsf{P}}\in{\mathcal{P}_{\textnormal{eff}}}}R_{1/\gamma}({\mathsf{Q}}\mid{\mathsf{P}})\right),$ (6.3)
and the supremum and infimum are attained by $X^{*}_{\gamma}\in{\mathcal{E}}$ and ${\mathsf{P}}^{*}_{\gamma}\in{\mathcal{P}_{\textnormal{eff}}}$ related by (6.2). Furthermore, $X^{*}_{\gamma}$ is the ${\mathsf{Q}}$-almost surely unique maximizer, and ${\mathsf{P}}^{*}_{\gamma}$ is the unique minimizer among the elements of ${\mathcal{P}_{\textnormal{eff}}}$ that are absolutely continuous with respect to ${\mathsf{Q}}$. Finally, $X^{*}_{\gamma}$ can be expressed in terms of ${\mathsf{P}}^{*}_{\gamma}$ by
@@
The equality in (6.3) and the form of the solutions were shown above. The ${\mathsf{Q}}$-almost sure uniqueness of $X^{*}_{\gamma}$ follows from strict concavity of the function $f(X)={\mathbb{E}}_{\mathsf{Q}}[U(X)]$ on ${\mathcal{E}}$. Since $\gamma>1$, the Rényi divergence is a decreasing transformation of the function $g({\mathsf{P}})={\mathbb{E}}_{\mathsf{Q}}[(d{\mathsf{P}}^{a}/d{\mathsf{Q}})^{1-1/\gamma}]$, which is thus maximized by ${\mathsf{P}}^{*}_{\gamma}$. The function $g({\mathsf{P}})$ is strictly concave on the set of elements of ${\mathcal{P}_{\textnormal{eff}}}$ that are absolutely continuous with respect to ${\mathsf{Q}}$. Thus ${\mathsf{P}}^{*}_{\gamma}$ is the unique maximizer of $g({\mathsf{P}})$, and the unique minimizer of the Rényi divergence. Finally, (6.4) is obtained by noting that
@@
Lardy et al. (2024, Section V) mention the e-variable $(d{\mathsf{P}}^{*}_{\gamma}/d{\mathsf{Q}})^{-1/\gamma}$ constructed using the reverse Rényi projection. This e-variable is strictly smaller than $X^{*}_{\gamma}$, because the denominator in (6.4) is strictly smaller than one by Jensen’s inequality.
@@
7 Universal inference is inadmissible